OpenStudy (lstrasmann):

Solve the equation: (3x - 2)(7x + 1) = 0. A. {2/3, -1/7} B. {-2/3, 1/7} C. 2/3 only D. -1/7 only E. no solution

OpenStudy (anonymous):

plug each value you into the equation to see which of those give you a true statement

OpenStudy (lstrasmann):

A and B have 2 values so how do I plug that in?? Like this? \[\left( 3\times2/3-2 \right)\left( 7\times-1/7+1 \right)\]

OpenStudy (espex):

Plug in one at a time. \[(3x-2)(7x+1)=0 \rightarrow (3(\frac{2}{3}) - 2)(7(\frac{2}{3}) + 1) = 0\] However, if you set each part equal to 0 you will have your answer: \[(3x-2)(7x+1)=0 \rightarrow 3x-2=0 \space, \space 7x+1=0\] and solve for x in each one.

OpenStudy (lstrasmann):

ok.. thank you!